OpenStudy (igreen):

Help? Let f(x) = -4x + 7 and g(x) = 2x - 6. Find \((g ° f)(1)\).

OpenStudy (igreen):

@hartnn

OpenStudy (igreen):

I don't know what the degree sign is supposed to mean..

hartnn (hartnn):

\((g ~o~f )(x)\) is read as "g of f of x" it is a composition of functions g and f

OpenStudy (igreen):

Oh, so I'm supposed to multiply the functions?

hartnn (hartnn):

to find \(g(f(1))\) you can also start this by directly finding f(1)

hartnn (hartnn):

f(1) is the value of f(x) when x =1

OpenStudy (igreen):

Isn't f(1) the same?

hartnn (hartnn):

f(x) = -4x + 7 plug in x =1 here you will get f(1)

OpenStudy (igreen):

f(x) = -4x + 7 f(x) = -4(1) + 7 f(x) = -4 + 7 f(x) = 3

OpenStudy (igreen):

Oh, now what do I do?

hartnn (hartnn):

f(x) = -4x + 7 f(1) = -4(1) + 7 f(1) = -4 + 7 f(1) = 3

hartnn (hartnn):

ok, now we have f(1) =3 \(\Large g(f(1))= g(3)\)

hartnn (hartnn):

so you just need g(3) to find g(3), plug in x = 3 in g(x)

OpenStudy (igreen):

Oh. \(g(x) = 2x - 6\) \(g(3) = 2(3) - 6\) \(g(3) = 6 - 6\) \(g(3) = 0\) Is that correct?

hartnn (hartnn):

\(\huge \checkmark\)

OpenStudy (igreen):

Awesome, didn't know it was that easy.

hartnn (hartnn):

:)

OpenStudy (conqueror):

How do you make that check? @hartnn

hartnn (hartnn):

type this `\(\huge \checkmark\)` that \checkmark is for the check \huge is to make the font very big

OpenStudy (conqueror):

\(\huge \checkmark)

OpenStudy (conqueror):

Oops, forgot the last slash

OpenStudy (conqueror):

\(\huge \checkmark\)

OpenStudy (conqueror):

Got it, thanks

hartnn (hartnn):

\(\Large Good ~ work ~ \ddot \smile \)